OpenStudy (anonymous):

write the equation of the line that passes through the point and is perpendicular to the given line ? (-12,6);y=4x-12

OpenStudy (dumbcow):

a perpendicular line will have a slope that is the opposite reciprocal to the original line. m = -1/4 use point-slope form y-6 = -1/4(x+12) y = -1/4x +3